40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 4 13:31:55 2012, Jenne, Update, General, The Plan 
    Reply  Tue Sep 4 13:44:17 2012, Manasa, Update, General, The Plan 
    Reply  Tue Sep 4 20:13:46 2012, Jenne, Update, General, The Plan - Tues evening version 
Message ID: 7340     Entry time: Tue Sep 4 20:13:46 2012     In reply to: 7335
Author: Jenne 
Type: Update 
Category: General 
Subject: The Plan - Tues evening version 

 Tues

* Hit center of ETMY, using input optics, PR3.

* Get IPANG out of vac, center QPD.

Wed

* AM: Riju do MC mode scans

* Starting right after 40m meeting, if not before: Lock MICH to get BS, ITMs aligned well

* Check if beam is hitting center of ITMX.

* Check for clipping around BS

     - Use Watek to look at beam at all 4 BS ports - make sure no clipping going into BS, after BS in the michelson, or the AS port

       - Use some old in-vac mirrors to direct beam out the BS door.  Cameras are waiting near BS chamber.

* Prepare glass beam dumps??

* IPPOS - make sure beam gets out of chamber

* Jan take photos of ETMX scattering setup

* Manasa take in-vac photos of all tables, for table layouts

* Jan / Manasa viewport transmission

* Install glass beam dumps?

* Install glass baffle at ETMY. Jan maybe install baffle at one of ITMs.

* ????

Thurs

* ????

* Check table levelling one last time on all tables. 

Fri

* Close all heavy doors.  (Access connector, ITMX, ITMY, BS, ETMX, ETMY? )

* Drag wipe test masses

* Start at ~10am?

Mon (if not Fri)

* Start pumping

ELOG V3.1.3-